OpenStudy (anonymous):

How many of each ticket were sold? children = x adult = y senior = z 800 tickets sold Formula: 800 = x+y+z Revenue is $3775, children tickets cost $3, adult $8, senior $5 Formula: 3775 = 3x+8y+5z Twice as many children tickets were sold than adults Formula: 2x=y So how many of each ticket were sold? PLEASE HELP! 800=x+y+z 3775=3x+8y+5z 2x=y

OpenStudy (anonymous):

pretty much i just needhelp finding x y and z

OpenStudy (tester97):

i think its 800=x+2x+z

OpenStudy (tester97):

since y=2x right?

OpenStudy (tester97):

then it would look like 3775= 3x+8(2x)+5z

OpenStudy (tester97):

then you would ditribute the 8 to 2x so it would look like 3775=3x+16x+5z

OpenStudy (tester97):

then you would put 3 for x and 5 for z then you would solve

OpenStudy (tester97):

let me know what you get when you are done @Ryan_Kochan

OpenStudy (anonymous):

after substititing y I have the two equations 800 = 3x + z and 3775 = 19x + 5z but thats as far as I can get because if I try to solve by elimination (multiply equation 1 by -5 and solve for x) then x = a negative number @tester97

OpenStudy (tester97):

ok now put 3 for x and 5 for z so it would look like this 800=3(3)+ (5) 3775=19(3)+5(5) now try to solve it now

OpenStudy (anonymous):

where did you get 3 for x and 5 for z from? @tester97

OpenStudy (tester97):

i got it from this part Revenue is $3775, children tickets cost $3, adult $8, senior $5 Formula: 3775 = 3x+8y+5z

OpenStudy (anonymous):

But x doesnt = 3, 3 is the cost of x. x is the amount of tickets of children sold

OpenStudy (anonymous):

help? @jim_thompson5910

OpenStudy (tester97):

did what i tell you make sense? @Ryan_Kochan

OpenStudy (anonymous):

No it doesnt make sense...i dont think that is right

OpenStudy (tester97):

well idk thats what my friend just told me and i typed exactly what she said sorry

OpenStudy (anonymous):

lol, its okay. thanks for trying though!

OpenStudy (tester97):

i always try to help others cos i always need help lol